# Unified Multi-Modal Interleaved Document Representation for Information Retrieval

### v1

@@ Abstract
Information Retrieval (IR) methods aim to identify relevant documents in response to a given query, which have gained remarkable attention due to their successful application in various natural language tasks. However, existing approaches typically consider only the textual information within the documents, which overlooks the fact that documents can contain multiple modalities, including texts, images, and tables. Further, they often segment each long document into multiple discrete passages for embedding, preventing them from capturing the overall document context and interactions between paragraphs. We argue that these two limitations lead to suboptimal document representations for retrieval. In this work, to address them, we aim to produce more comprehensive and nuanced document representations by holistically embedding documents interleaved with different modalities. Specifically, we achieve this by leveraging the capability of recent vision-language models that enable the processing and integration of text, images, and tables into a unified format and representation. Moreover, to mitigate the information loss from segmenting documents into passages, instead of representing and retrieving passages individually, we further merge the representations of segmented passages into one single document representation, while we additionally introduce a reranking strategy to decouple and identify the relevant passage within the document if necessary. Then, through extensive experiments on diverse information retrieval scenarios considering both the textual and multimodal queries, we show that our approach substantially outperforms relevant baselines, thanks to the consideration of the multimodal information interleaved within the documents in a unified way.
@@ 1 Introduction
Information Retrieval (IR) is the task of fetching relevant documents from a large corpus in response to an input query, which becomes a fundamental process to various real-world applications including web search engines and question-answering systems (Shah et al., 2019; Lewis et al., 2020; Guu et al., 2020). Specifically, to retrieve documents for the query, traditional approaches have focused on their textual representations, utilizing either sparse retrieval methods such as TF-IDF and BM25 (Robertson et al., 1994; Jones, 2004), which rely on exact term matching between the query and document, or dense retrieval methods such as DPR and ANCE (Karpukhin et al., 2020; Xiong et al., 2021), which leverage neural embeddings of the query and document text to capture semantic similarities between them over a continuous vector space. Recently, dense retrieval methods have gained more popularity over sparse methods due to their capability to capture semantic nuances and context beyond simple keyword matching, leading to multiple successes with improved performance.
Despite their huge successes, existing (dense) retrieval methods face a couple of severe challenges. First, they primarily rely on the textual data for document embedding and retrieval, overlooking the fact that modern documents often contain multimodal content, such as images and tables (beyond the plain text), which can carry critical information that may be essential for accurately understanding and retrieving the relevant documents. To be specific, a diagram within a medical article can more effectively represent the structure of a molecule or the progression of a disease, offering more clarity that would be difficult to achieve with text alone, and omitting such multimodal content can lead to an incomplete understanding (and potentially inaccurate retrieval) of the documents. Also, the segmentation of long documents into discrete passages, which is commonly employed by retrieval models to handle the length limitation for embeddings, may prevent models from capturing the full context and the intricate relationships between different parts of the document, ultimately leading to suboptimal retrieval performance. It is worthwhile noting that, concurrent to our work, while there has been recent work that screen captures the document and then embed its screenshots (to consider different modalities in a unified format) (Faysse et al., 2024; Ma et al., 2024a), not only its content (such as paragraphs, images, and tables) can be fragmented into different sub-images, leading to the loss of contextual coherence across the entire document, but also the visual representation of text may hinder the model’s ability to capture the semantic relationships present in the original textual data, while increasing the image resolution leads to the critical concern on the memory requirements.
In this work, we introduce a novel approach to holistic document embedding for IR, which addresses the aforementioned challenges by representing and retrieving the documents interleaved with different modalities in a unified manner. Specifically, our method revolves around the recent advance of Vision-Language Models (VLMs), which enables the processing and integration of multimodal content (such as text, images, and tables) directly into a single token sequence, thereby preserving the context and relationships between various parts of the document, unlike the previous approaches that rely on the fragmented visual representations. Furthermore, in cases where the number of tokens in a document is large and exceeds the capacity of a single context window of VLMs, we propose a strategy to segment the document into manageable passages, each represented within the token limit, and combine these passage representations into a unified document representation, which differs from existing IR approaches that independently represent and retrieve at the passage level, potentially losing the overall document context. Lastly, to accurately identify only the relevant sections within the lengthy documents, we introduce a reranking mechanism that is trained to pinpoint the passage most pertinent to the query (among all the other passages within the document), effectively allowing for both the coarse-grained document-level matching and the fine-grained passage-level retrieval. We provide the visual illustrations of the overall pipeline of IDentIfy against prior work in Figure 1. We refer to our overall method as Interleaved Document Information Retrieval System (IDentIfy).
We experimentally validate the effectiveness of IDentIfy on four different benchmark datasets, considering both the text-only and multimodal queries. On a battery of tests conducted, we then observe that our approach substantially outperforms relevant baselines that consider only the uni-modality for document representations, thanks to the consideration of multimodal content. Further, we find that the strategy to represent the whole document with its single representation (by merging embeddings of its splits if necessary) is superior to the approach of individually representing them for document retrieval, but also performing reranking over the sections of the retrieved document is superior to the approach of directly retrieving those sections, which confirm the efficacy of the proposed retrieval and reranking pipeline for document and passage retrieval, respectively.
@@
Figure 1: Comparison of different IR approaches. (a): Conventional methods use a small portion of the text within the document for its representation. (b): Recent methods use first-page screenshot images to represent the document. (c): Our approach leverages the full contextual information within documents interleaved with multiple modalities by considering them in their original format, and is capable of pinpointing relevant sections.
@@ 2 Related Work
Information Retrieval (IR) is the task of accurately finding documents relevant to a given query from a large corpus, such as Wikipedia, which has been a crucial component for a variety of applications, including search engines, question-answering systems, and conversational agents (Zhu et al., 2023; Gao et al., 2023; Ram et al., 2023; Shi et al., 2024; Jeong et al., 2024a). Specifically, to retrieve the relevant documents, earlier IR approaches measured similarity between queries and documents based on their lexical term matching, such as BM25 and TF-IDF (Robertson et al., 1994; Jones, 2004). Yet, these methods often struggled to capture the semantic nuances beyond surface-level term overlaps. To overcome this, along with advancements in language models (Devlin et al., 2019; Liu et al., 2019). there has been dense retrieval approaches that embed both the queries and documents into a shared dense vector space (Karpukhin et al., 2020; Xiong et al., 2021), enabling the calculation of semantic similarity between them more effectively by capturing the deeper contextual information. However, previous IR studies have mainly focused on enhancing the textual representations of queries and documents, while overlooking the fact that documents often consist of diverse modalities (such as images and tables) beyond text, which can potentially provide richer context and aid in more accurate retrieval (Liu et al., 2021; Jeong et al., 2024b).
@@
Recent studies in IR have expanded the focus from purely text-based retrieval models to those that consider other modalities, such as images (Radford et al., 2021; Xiao et al., 2024), tables (Herzig et al., 2021; Chen et al., 2024) and graphs (Baek et al., 2023); however, the majority of these approaches (Zhou et al., 2024; Long et al., 2024; Lerner et al., 2024; Nowak et al., 2024; Caffagni et al., 2024) have primarily explored how to process the multimodal queries, meanwhile, they often overlook the equally important multimodal characteristics of the documents being retrieved. Specifically, we argue that, while incorporating multimodal elements in queries has expanded the range and diversity of query types that IR systems can handle, considering the multimodal nature of the documents can lead to a more holistic representation of retrieval targets, which can ultimately lead to enhancing the overall retrieval performance. In efforts to handle diverse multimodal elements within documents, there are concurrent studies that have proposed to capture screenshots of documents, such as PDFs (Faysse et al., 2024) or Wikipedia web pages (Ma et al., 2024a), and subsequently encoding them through vision models (Ding et al., 2024). However, these methods are not only limited by factors, such as image resolution and computational memory, constraining their application to documents longer than a single page [^1] For instance, Ma et al. (2024a) requires processing 9.8k image tokens just to process a single-page document, and it results in 2TB of memory for handling the entire Wikipedia corpus, which is not much practical., but also fall short by treating the diverse modalities within a document as a single visual entity, leading to suboptimal document representations that fail to effectively capture the nuanced interdependence between text and images. Furthermore, they do not address the critical issue of splitting documents into smaller fragments (e.g., sub-images), which may disrupt the holistic contextual view of the entire document.
@@
Recent Vision-Language Models (VLMs) have emerged as a powerful tool for jointly processing visual and textual data, combining the image understanding capabilities of visual encoders (Radford et al., 2021; Zhai et al., 2023) with the advanced reasoning abilities of language models (OpenAI, 2022; 2023a). These models have achieved remarkable performance across diverse vision-language (VL) tasks (such as image captioning and visual question answering) (Dai et al., 2023; OpenAI, 2023b), with the substantially limited attention on their applications to IR. We note that the latest developments in this field have particularly focused on enabling VLMs to handle interleaved, multimodal content, which involves a mixed sequence of images and text (Zhang et al., 2023; Li et al., 2024). In particular, LLaVA-NeXT-Interleave (Li et al., 2024) introduces a fine-tuning approach that specifically enhances the VLMs’ capacity to understand complex interleavings of multiple images and text within a single context. Drawing inspiration from these advances, in this work, we propose to harness the capabilities of VLMs to create unified embeddings for documents interleaved with text and images (as well as tables) for IR, which is a big shift from even the recent IR approach (Ma et al., 2024b) that still embeds the documents with the recent but text-based models like Llama (Touvron et al., 2023a; b), failing to fully capture the diverse multimodal content.
@@ 3.1 Preliminary
[…]
@@
We now turn to describing Vision-Language Models (VLMs), which are designed to jointly encode the textual and visual information in a unified token framework. We note that these models are generally comprised of two main components: a visual encoder and a language model, interconnected through a projection layer. Specifically, given an input document that may contain interleaved modalities (e.g., text and images), the visual encoder extracts high-level visual features from (multiple) images embedded within the document, mapping them into a latent space. Then, these visual features are transformed into a sequence of visual tokens via the projection layer, represented as follows: $\mathbf{V}\!\in\!\mathbb{R}^{V\!\times\!d_{\text{emb}}}$ where $V$ denotes the visual token length and $d_{\text{emb}}$ is the token dimension size. Similarly, for the textual content embedded within the document alongside images, the language model uses a word embedding layer to convert the input text into a sequence of text tokens, represented as follows: $\mathbf{L}\!\in\!\mathbb{R}^{L\!\times\!d_{\text{emb}}}$ where $L$ denotes the token length of text.
In this work, we also propose to account for tables that are an integral modality for holistically representing the full content of documents. However, in contrast to text and images that have dedicated processing layers within the VLM architectures, tables do not have a specific representation layer. Nevertheless, we argue that recent VLMs are pre-trained on diverse web data, and subsequently they are implicitly learned to handle the table structures formatted in HTML. Consequently, we treat HTML-format table data as a linearized sequence of HTML words, applying the same word embedding layer as is used for plain text. To be formal, this process converts the table content into table tokens, as follows: $\mathbf{T}\!\in\mathbb{R}^{T\!\times\!d_{\text{emb}}}$ where $T$ is the token length of the table. Lastly, once extracted, the visual tokens, text tokens, and table tokens are concatenated (to form a unified token sequence) and then passed through the remaining layers of VLMs, to capture both uni- and cross-modal relationships across different modalities, enabling the comprehensive understanding of the input document.
@@ 3.2 Retriever
We now turn to explaining how we design a retriever specifically tailored for multimodal interleaved document retrieval. In particular, to effectively retrieve documents that contain multiple modalities, our approach leverages a VLM capable of processing text, images, and tables within a single document. Further, following the standard practice of existing retrieval architectures (Karpukhin et al., 2020; Xiong et al., 2021), we use a dual-encoder structure, which consists of a query encoder and a section encoder, both are based on the VLM, which is illustrated in Figure 2 (a).
@@
The remaining step to discuss here is how to train those two query and document retrievers for IR. Recall that the goal of the retriever is to assess a relevance score between the query and the document. To achieve this goal, we use a contrastive learning loss based upon the query and document representations, whose objective is to assign higher similarity scores to relevant documents (positive samples) and lower scores to irrelevant ones (negative samples) for the query, formulated as follows:
$\mathcal{L}_{\text{retriever}}\!$ $=\!-\frac{1}{B}\sum^{B}_{i=1}\log\left(\frac{\texttt{sim}(\mathbf{Z}_{\mathrm{Q}_{i}},\mathbf{Z}_{\mathrm{D}_{i}})}{\sum^{B}_{j=1}\texttt{sim}(\mathbf{Z}_{\mathrm{Q}_{i}},\mathbf{Z}_{\mathrm{D}_{j}})}\right),\;\;\texttt{sim}\left(\mathbf{Z}_{\mathrm{Q}},\;\mathbf{Z}_{\mathrm{D}}\right)\!=\!\frac{\mathbf{Z}_{\mathrm{Q}}^{\top}\mathbf{Z}_{\mathrm{D}}}{\lVert\mathbf{Z}_{\mathrm{Q}}\rVert\lVert\mathbf{Z}_{\mathrm{D}}\rVert},$

### v2

# Unified Multimodal Interleaved Document Representation for Retrieval

@@ Abstract
Information Retrieval (IR) methods aim to identify documents relevant to a query, which have been widely applied in various natural language tasks. However, existing approaches typically consider only the textual content within documents, overlooking the fact that documents can contain multiple modalities, including images and tables. Also, they often segment each long document into multiple discrete passages for embedding, which prevents them from capturing the overall document context and interactions between paragraphs. To address these two challenges, we propose a method that holistically embeds documents interleaved with multiple modalities by leveraging the capability of recent vision-language models that enable the processing and integration of text, images, and tables into a unified format and representation. Moreover, to mitigate the information loss from segmenting documents into passages, instead of representing and retrieving passages individually, we further merge the representations of segmented passages into one single document representation, while we additionally introduce a reranking strategy to decouple and identify the relevant passage within the document if necessary. Then, through extensive experiments on diverse IR scenarios considering both the textual and multimodal queries, we show that our approach substantially outperforms relevant baselines, thanks to the consideration of the multimodal information within documents.
Unified Multimodal Interleaved Document Representation for Retrieval
Jaewoo Lee1,2∗   Joonho Ko1∗   Jinheon Baek1∗   Soyeong Jeong1   Sung Ju Hwang1,2 KAIST1 DeepAuto2 jwlee8877@gmail.com, {joonho.ko, jinheon.baek, starsuzi}@kaist.ac.kr
**footnotetext: Equal contribution
@@ 1 Introduction
Information Retrieval (IR) is the task of fetching relevant documents from a large corpus in response to an input query, which plays a critical role in various real-world applications including web search engines and question-answering systems (Shah et al., 2019; Lewis et al., 2020; Guu et al., 2020). Over the years, IR methods have evolved significantly, with approaches broadly categorized into sparse and dense retrieval paradigms. Specifically, sparse retrieval methods Robertson et al. (1994); Jones (2004) focus on lexical overlap between queries and documents; meanwhile, dense retrieval methods Karpukhin et al. (2020); Xiong et al. (2021) utilize neural embeddings to represent queries and documents in a continuous vector space. Note that, recently, dense retrieval methods have gained more popularity over sparse methods due to their capability to capture semantic nuances and context beyond simple keyword matching, leading to multiple successes with improved performance.
@@
Figure 1: Comparison of different IR approaches. (a): Conventional methods use a small portion of the text within the document for its representation. (b): Recent methods use first-page screenshot images to represent the document. (c): Our approach leverages the full contextual information within documents interleaved with multiple modalities by considering them in their original format, and is further capable of pinpointing relevant sections for the query.
Despite their huge successes, existing (dense) retrieval methods face a couple of severe challenges. First, they primarily rely on the textual data for document embedding and retrieval, overlooking the fact that modern documents often contain multimodal content, such as images and tables (beyond the plain text), which can carry critical information that may be essential for accurately understanding and retrieving the relevant documents. To be specific, a diagram within a medical article can more effectively represent the structure of a molecule or the progression of a disease, offering more clarity that would be difficult to achieve with text alone, and omitting such multimodal content can lead to an incomplete understanding (and potentially inaccurate retrieval) of the documents. Also, the segmentation of long documents into discrete passages, which is commonly employed by existing retrieval models to handle the length limitation for embeddings, may prevent models from capturing the full context and the intricate relationships between different parts of the document, ultimately leading to suboptimal retrieval performance. It is worthwhile noting that, concurrent to our work, while there has been recent work that screen captures the document and then embed its screenshots (to consider different modalities in a unified format) (Faysse et al., 2024; Ma et al., 2024), not only its content (such as paragraphs, images, and tables) can be fragmented into different sub-images, leading to the loss of contextual coherence across the entire document, but also the visual representation of text may hinder the model’s ability to capture the semantic relationships present in the original textual data, while increasing the image resolution leads to the concern on the memory requirements.
In this work, we introduce a novel approach to holistically represent documents for IR, which addresses the aforementioned challenges by representing and retrieving the documents interleaved with multiple modalities in a unified manner (See Figure 1). Specifically, our method revolves around the recent advance of Vision-Language Models (VLMs), which enable the processing and integration of multimodal content (such as text, images, and tables) directly into a single token sequence, thereby preserving the context and relationships between various parts of the document, unlike the previous approaches that rely on the fragmented visual representations. Also, in cases where the number of tokens in a document is large and exceeds the capacity of a single context window of VLMs, we propose a strategy to segment the document into passages, each represented within the token limit, and combine these passage embeddings into a unified document representation. This strategy differs from existing IR approaches that independently represent and retrieve at the passage level, potentially losing the overall document context. Lastly, to accurately identify only the relevant sections within the retrieved lengthy document, we introduce a reranking mechanism that is trained to pinpoint the passage most pertinent to the query (among all the other passages within the document), allowing for both the coarse-grained document-level matching and fine-grained passage-level retrieval. We refer to our overall framework as Interleaved Document Information Retrieval System (IDentIfy).
We experimentally validate the effectiveness of IDentIfy on four benchmark datasets, considering both the text-only and multimodal queries. On a battery of tests conducted, we observe that our approach substantially outperforms relevant baselines that consider only the uni-modality for document representations, thanks to the consideration of multimodal content. Further, we find that the strategy to represent the whole document with its single representation (by merging embeddings of its splits) is superior to the approach of individually representing them for document retrieval, but also performing reranking over the sections of the retrieved document is superior to the approach of directly retrieving those sections, which confirm the efficacy of the proposed retrieval and reranking pipeline for document and passage retrieval, respectively.
@@ 2 Related Work
Information Retrieval (IR) involves finding documents relevant to a query, which plays a crucial role in applications such as search and question-answering (Zhu et al., 2023; Gao et al., 2023; Ram et al., 2023; Shi et al., 2024; Jeong et al., 2024a). Earlier IR approaches measured the similarity between queries and documents based on their lexical term matching, such as BM25 and TF-IDF (Robertson et al., 1994; Jones, 2004). Yet, these methods often struggled to capture the semantic nuances beyond surface-level term overlaps. To overcome this, along with advancements in language models (Devlin et al., 2019; Liu et al., 2019), there have been dense retrieval approaches that embed both the queries and documents into a shared dense vector space (Karpukhin et al., 2020; Xiong et al., 2021), enabling the calculation of semantic similarity between them more effectively by capturing the deeper contextual information. Yet, previous studies have mainly focused on enhancing the textual representations of queries and documents, while overlooking the multimodal nature of documents beyond text, which can potentially provide richer context and aid in more accurate retrieval (Liu et al., 2021; Jeong et al., 2024b).
@@
Recent studies in IR have expanded the focus from purely text-based retrieval models to those that consider other modalities, such as images (Radford et al., 2021; Xiao et al., 2024), tables (Herzig et al., 2021; Chen et al., 2024) and graphs (Baek et al., 2023); however, the majority of these approaches (Zhou et al., 2024; Long et al., 2024; Lerner et al., 2024; Nowak et al., 2024; Caffagni et al., 2024) have primarily explored how to process the multimodal queries, meanwhile, they often overlook the equally important multimodal characteristics of the documents being retrieved. In efforts to handle diverse multimodal elements within documents, there are concurrent studies that have proposed to capture screenshots of documents, such as PDFs (Faysse et al., 2024; Cho et al., 2024) or Wikipedia web pages (Ma et al., 2024), and subsequently encoding them through vision models (Ding et al., 2024). However, these methods are not only limited by factors, such as image resolution and computational memory, constraining their application to documents longer than a single page [^1] For instance, Ma et al. (2024) requires processing 9.8k image tokens just to process a single-page document, and it results in 2TB of storage for handling the entire Wikipedia corpus, which may not be practical., but also fall short by treating the diverse modalities within a document as a single visual entity, leading to suboptimal document representations that may fail to effectively capture the nuanced interdependence between text and images. Also, while there are concurrent studies Jiang et al. (2024b); Lin et al. (2024) that consider images and text as retrieval targets, they primarily focus on representing image-text pairs and their retrieval, rather than addressing the holistic representation of documents that include multiple images and another modality (tables). Finally, all the aforementioned work does not address the issue of splitting documents into smaller fragments (passages or sub-images), which may disrupt the holistic contextual view of the entire document.
@@
Recently developed Vision-Language Models (VLMs) have emerged as a powerful tool for jointly processing visual and textual data, which combine the image understanding capabilities of visual encoders (Radford et al., 2021; Zhai et al., 2023) with the advanced reasoning abilities of language models (OpenAI, 2022, 2023a). These models have achieved remarkable performance across diverse vision-language (VL) tasks (such as image captioning and visual question answering) (Dai et al., 2023; OpenAI, 2023b), with the substantially limited attention on their applications to IR. We note that the latest developments in this field have particularly focused on enabling VLMs to handle interleaved, multimodal content, which involves a mixed sequence of images and text (Zhang et al., 2023; Li et al., 2024b). In particular, LLaVA-NeXT-Interleave (Li et al., 2024b) introduces a fine-tuning approach that specifically enhances the VLMs’ capacity to understand complex interleavings of multiple images and text within a single context. Drawing inspiration from these advances, we propose to harness the capabilities of VLMs to create unified embeddings for documents interleaved with text and images (as well as tables).
@@ 3.1 Preliminaries
IR is formally defined as the task of identifying a set of relevant documents $\{{\bm{d}}_{1},{\bm{d}}_{2},\ldots,{\bm{d}}_{k}\}\subseteq\mathcal{D}$ from a large corpus $\mathcal{D}$, given an input query ${\bm{q}}$. Here, each query ${\bm{q}}$ and document ${\bm{d}}$ are represented as a sequence of tokens: ${\bm{q}}=[q_{1},q_{2},\ldots,q_{n}]$ and ${\bm{d}}=[d_{1},d_{2},\ldots,d_{m}]$, and traditional IR approaches typically consider these tokens as purely textual elements. However, we propose to extend this assumption to have the tokens of both the textual and visual content, to capture the multimodal nature of many real-world documents. Then, this new extension raises important questions of how can both the textual and visual content be represented within a unified token framework, and how can these multimodal tokens be seamlessly integrated and encoded for document representations. To answer them, we harness the power of recent vision-language models below.
Figure 2: Overview of the proposed IDentIfy. (a): In our document retriever, a query encoder represents a query (purple), and sections are encoded with a section encoder whose embeddings averaged to form a document representation (blue). Contrastive learning loss (red) is used for training the document retriever. (b): Reranker scores query-section relevance with the concatenation of the query and section, trained using Binary Cross-Entropy loss.
@@
We now turn to describing Vision-Language Models (VLMs), which are designed to jointly encode the textual and visual information in a unified token framework. These models are generally comprised of two main components: a visual encoder and a language model, interconnected through a projection layer. Specifically, given the document that may contain interleaved modalities (e.g., text and images), the visual encoder extracts high-level visual features from (multiple) images embedded within the document, mapping them into a latent space. Then, these visual features are transformed into a sequence of visual tokens via the projection layer, represented as follows: $\mathbf{V}\!\in\!\mathbb{R}^{V\!\times\!d_{\text{emb}}}$, where $V$ denotes the visual token length and $d_{\text{emb}}$ is the token dimension size. Similarly, for the textual content embedded within the document, the language model uses a word embedding layer to convert the input text into a sequence of tokens, as follows: $\mathbf{L}\!\in\!\mathbb{R}^{L\!\times\!d_{\text{emb}}}$, where $L$ denotes the token length of text.
In this work, we also propose to account for tables that are an integral modality for holistically representing the full content of documents. However, in contrast to text and images that have dedicated processing layers within the VLM architectures, tables do not have a specific representation layer. Nevertheless, we argue that recent VLMs are pre-trained on diverse web data, and subsequently they are implicitly learned to handle the table structures formatted in HTML. Consequently, we treat HTML-format table data as a linearized sequence of HTML words, applying the same word embedding layer as is used for plain text. To be formal, this process converts the table content into table tokens, as follows: $\mathbf{T}\!\in\mathbb{R}^{T\!\times\!d_{\text{emb}}}$, where $T$ is the token length of the table. Lastly, once extracted, the visual tokens, text tokens, and table tokens are concatenated (to form a unified token sequence) and then passed through the remaining layers of VLMs, to capture both uni- and cross-modal relationships across different modalities, ultimately enabling the comprehensive understanding of the documents.
@@ 3.2 Retriever
We now explain how we design a retriever specifically tailored for multimodal interleaved document retrieval. In particular, our approach leverages a VLM capable of processing text, images, and tables within a single document. Further, following the standard practice of existing retrieval architectures (Karpukhin et al., 2020; Xiong et al., 2021), we use a dual-encoder structure, which consists of a query encoder and a section encoder, both are based on the VLM, illustrated in Figure 2 (a).
@@
The remaining step to discuss here is how to train those two query and section encoders for IR. Recall that the goal of the retriever is to assess a relevance score between the query and the document. To achieve this goal, we use a contrastive learning loss based upon the query and document representations, whose objective is to assign higher similarity scores to relevant documents (positive samples) and lower scores to irrelevant ones (negative samples) for the query, formulated as follows:
$\mathcal{L}_{\text{retriever}}\!$ $=\!-\frac{1}{B}\sum^{B}_{i=1}\log\left(\frac{\phi\left(\mathbf{Z}_{\mathrm{Q}_{i}},\mathbf{Z}_{\mathrm{D}_{i}}\right)}{\sum^{B}_{j=1}\phi\left(\mathbf{Z}_{\mathrm{Q}_{i}},\mathbf{Z}_{\mathrm{D}_{j}}\right)}\right),$
$\phi\left(\mathbf{a},\;\mathbf{b}\right)\!$ $=\!\exp\left(\frac{\mathbf{a}^{\top}\mathbf{b}}{\lVert\mathbf{a}\rVert\lVert\mathbf{b}\rVert}\right),$ (1)